# Finite difference method; stability; quantitative finance; stochastic volatility

### 03 Jan 2024 v1

@@ 2 Quick summary of the SWIFT method
[…]
$\phi_{m,k}(x)$ $=2^{\frac{m}{2}}\phi\left(2^{m}x-k\right)\,,$ (4)
$\phi(x)$ $=\mathrm{sinc}(x)=\frac{\sin\pi x}{\pi x}$ (5)
@@
$V_{m,k}$ $=K\frac{2^{\frac{m}{2}}}{J}\sum_{j=1}^{J}\Re\left\{e^{-i\omega_{j}k}\int_{0}^{c}(1-e^{y})e^{i\omega_{j}2^{m}y}\mathop{}\!\kern 0.0pt\mathrm{d}y\right\}\,.$ (8)
Le Floc’h (2020) proposes another formula, centered on the forward, such that the characteristic function is only evaluated in the center ($x=0$):
@@ 3 Choice of SWIFT parameters
and $c$ itself is updated according to the total density criteria (Equation 13) while $\kappa$ and $J$ are recomputed from the new $c$ in the iterative procedure. The estimate of $m$ is still somewhat critical to be able to reach the accuracy set for $\epsilon_{f}$: if $\epsilon_{m}$ is set to be too large, $m$ may be too small to reach a the accuracy set for $\epsilon_{f}$. This may explain why Romo and Ortiz-Gracia (2021) proposed to circumvent $\epsilon_{m}$ altogether (at the cost of a strong sensitivity on the choice of $L$ as explained above). In practice, we found the estimate of $m$ to be too conservative when using similar $\epsilon_{m}$ and $\epsilon_{f}$ thresholds, so the procedure often works, but may lead to a 2x or 4x increase of the number of points used. The range $[-c,c]$ is adjusted using the minimum and maximum quoted log-moneyness to $[x_{\min}-c,x_{\max}+c]$. In this note, we don’t apply this adjustment but use the payoff formula centered on the forward of Le Floc’h (2020). The interval adjustment along with the strike based formula would not change any of the observations made in this note. The choice $\log_{2}(J)=\lceil\log_{2}(\pi\kappa)\rceil$ is motivated by the range allowed in the series which constitutes the upper bound of the error in the sinc function approximation in (Ortiz-Gracia and Oosterlee, 2016). In reality, the more direct upper bound from the Euler-Maclaurin expansion (Equation 9) is much stricter and the choice $\log_{2}(J)=\lceil\log_{2}(\kappa)\rceil$ is enough for the FFT algorithm on $N=2J$ to generate the indices in the interval $\{-\kappa+1,...,\kappa\}$.
@@ 6 Conclusion
There are not many improvements to propose on the SWIFT method as described in (Maree et al., 2017; Leitao et al., 2018): more precise approximations of sinc do not translate into a more precise option price. The initialization procedure of Leitao et al. (2018) is more agnostic to the choice of initial range. And in this regard, the SWIFT method is more robust than the COS method.

### 07 Jan 2024 v3

@@ 2 Quick summary of the SWIFT method
$\phi_{m,k}(x)=2^{\frac{m}{2}}\phi\left(2^{m}x-k\right)$ $\,,\quad\phi(x)=\mathrm{sinc}(x)=\frac{\sin\pi x}{\pi x}\,,$ (3)
@@
$V_{m,k}$ $=K\frac{2^{\frac{m}{2}}}{J}\sum_{j=1}^{J}\Re\left\{e^{-i\omega_{j}k}\int_{0}^{c}(1-e^{y})e^{i\omega_{j}2^{m}y}\mathop{}\!\kern 0.0pt\mathrm{d}y\right\}\,,$ (6)
where $c=\frac{\kappa}{2^{m}}$. Le Floc’h (2020) proposes another formula, centered on the forward, such that the characteristic function is only evaluated in the center ($x=0$):
@@ 3 Choice of SWIFT parameters
and $c$ itself is updated according to the total density criteria (Equation 11) while $\kappa$ and $J$ are recomputed from the new $c$ in the iterative procedure. This is more in line with (Ortiz-Gracia and Oosterlee, 2016), where $c=\frac{\kappa-1}{2^{m}}$ is used in the calculation of the put payoff coefficient. The estimate of $m$ is still somewhat critical to be able to reach the accuracy set for $\epsilon_{f}$: if $\epsilon_{m}$ is set to be too large, $m$ may be too small to reach a the accuracy set for $\epsilon_{f}$. This may explain why Romo and Ortiz-Gracia (2021) proposed to circumvent $\epsilon_{m}$ altogether (at the cost of a strong sensitivity on the choice of $L$ as explained above). In practice, we found the estimate of $m$ to be too conservative when using similar $\epsilon_{m}$ and $\epsilon_{f}$ thresholds, so the procedure often works, but may lead to a 2x or 4x increase of the number of points used. The range $[-c,c]$ is adjusted using the minimum and maximum quoted log-moneyness to $[x_{\min}-c,x_{\max}+c]$. In this note, we don’t apply this adjustment but use the payoff formula centered on the forward of Le Floc’h (2020). The interval adjustment along with the strike based formula would not change any of the observations made in this note. The choice $\log_{2}(J)=\lceil\log_{2}(\pi\kappa)\rceil$ is motivated by the range allowed in the series which constitutes the upper bound of the error in the sinc function approximation in (Ortiz-Gracia and Oosterlee, 2016). In reality, the more direct upper bound from the Euler-Maclaurin expansion (Equation 7) is much stricter and the choice $\log_{2}(J)=\lceil\log_{2}(\kappa)\rceil$ is enough for the FFT algorithm on $N=2J$ to generate the indices in the interval $\{-\kappa+1,...,\kappa\}$.
@@ 6 Conclusion
There are not many improvements to propose on the SWIFT method as described in (Maree et al., 2017; Leitao et al., 2018): more precise approximations of sinc do not translate into a more precise option price. The initialization procedure of Leitao et al. (2018) is more agnostic to the choice of initial range. And in this regard, the SWIFT method is more robust than the COS method. We also proved that the expansions were converging with a slightly smaller $J$ than the choice presented in (Ortiz-Gracia and Oosterlee, 2016) and subsequent papers.